OpenStudy (anonymous):

find two numbers whose sum is 13 and the sum of whose square is 145

OpenStudy (bahrom7893):

x+y=13 x^2+y^2=145

OpenStudy (bahrom7893):

x = 13-y (13-y)^2+y^2=145

OpenStudy (bahrom7893):

solve the 2nd equation: 169-26y+y^2+y^2=145 2y^2 -26y + 169 - 145 = 0

OpenStudy (bahrom7893):

2y^2 - 26y + 24 = 0 y^2 - 13y + 12 = 0

OpenStudy (bahrom7893):

D = 169 - 48 = 121 y1 = (13 - 11)/2 = 1 y2 = (13+11)/2 = 12 So now you have two y values..

OpenStudy (bahrom7893):

Find the two x values: x1 = 13-y1 = 13-1 = 12 x2 = 13-y2 = 13 -12 = 1

OpenStudy (bahrom7893):

So apparently the two numbers u need are 12 and 1..

myininaya (myininaya):

very good bahrom

OpenStudy (anonymous):

did you have the same answer??...myininaya??

myininaya (myininaya):

yes his answers are glamorous

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

Heres another solution that uses the symmetry of the problem:

myininaya (myininaya):

very nice joe

OpenStudy (anonymous):

tnx!!